honestyprewitt:

Given the trinomial, what is the value of the coefficient 'B' in the factored form? 2x2 − 12xy − 32y2 = 2(x − 8y)(x + By) B= ____________

jhonyy9:

- like a first step factoriz out the 2 and will get 2x^2 -12xy -32y^2 = 2(x^2 -6xy -16y^2) = 2(x-8y)(x+2y) so B=2 hope helped